# Towards Universal Soccer Video Understanding

### v1

@@ A.1 Dataset Format
"timestamp": "2022-08-07 21:00:00", # Match start time
"score": "1 - 2", # Final score
"home_team": "Manchester Utd", # Home team name
"away_team": "Brighton", # Away team name
"home_formation": "4 - 3 - 3", # Home team formation
"away_formation": "3 - 4 - 2 - 1", # Away team formation
"venue": "Old Trafford (Manchester)", # Venue and city
"capacity": "75 635", # Stadium capacity
"attendance": "73 711", # Number of attendees
@@
"country": "Eng", # Referee’s nationality
"name": "Paul Tierney" # Referee’s name
@@
"players_name": "Caicedo M.", # Player’s abbreviated name
"players_number": "25", # Jersey number
"Full Name": "Moises Caicedo", # Player’s full name
"players_rating": 7.6, # Post-match rating
"Country": "Ecuador", # Player’s nationality
"Image URL": "https://static.flashsc...", # URL to player image
"Role": "Midfielder", # On-field role
"Age and Birthdate": "22, (02.11.2001)", # Age and birth date
"Market Value": "€89.4m" # Player’s market value
@@
"half": 1, # Match half (1 or 2)
"time_stamp": "00:16", # Timestamp within the half
"comments_type": "shot off target", # Event type
"comments_text": "A mistake by Leandro Trossard (Brighton)...", # Commentary text
"comments_text_anonymized": "A mistake by [PLAYER]([TEAM])..." # Commentary after anonymization

### v2

@@ A.1 Dataset Format
{ "timestamp": "2022-08-07 21:00:00", # Match start time "score": "1 - 2", # Final score "home_team": "Manchester Utd", # Home team name "away_team": "Brighton", # Away team name "home_formation": "4 - 3 - 3", # Home team formation "away_formation": "3 - 4 - 2 - 1", # Away team formation "venue": "Old Trafford (Manchester)", # Venue and city "capacity": "75 635", # Stadium capacity "attendance": "73 711", # Number of attendees }
@@
{ "country": "Eng", # Referee’s nationality "name": "Paul Tierney" # Referee’s name }
@@
{ "players_name": "Caicedo M.", # Player’s abbreviated name "players_number": "25", # Jersey number "Full Name": "Moises Caicedo", # Player’s full name "players_rating": 7.6, # Post-match rating "Country": "Ecuador", # Player’s nationality "Image URL": "https://static.flashsc...", # URL to player image "Role": "Midfielder", # On-field role "Age and Birthdate": "22, (02.11.2001)", # Age and birth date "Market Value": "€89.4m" # Player’s market value }
@@
{ "half": 1, # Match half (1 or 2) "time_stamp": "00:16", # Timestamp within the half "comments_type": "shot off target", # Event type "comments_text": "A mistake by Leandro Trossard (Brighton)...", # Commentary text "comments_text_anonymized": "A mistake by [PLAYER]([TEAM])..." # Commentary after anonymization }